OpenStudy (anonymous):

Solve 3x^2 + 2 = 11 - x^2 A. {±3/2} B. {±√3} C. {}

geerky42 (geerky42):

Add x^2 to both sides

OpenStudy (anonymous):

I think it's A, am i correct?

OpenStudy (anonymous):

did you solve it?

OpenStudy (anonymous):

I'm working on it

OpenStudy (anonymous):

tell me when you are sure :P

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I mean i think it's B

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

its not B

OpenStudy (anonymous):

ofcrs its A

OpenStudy (vishweshshrimali5):

Its A.

OpenStudy (anonymous):

4x^2 = 9 x^2 =9/4 x= +- 3/2

OpenStudy (anonymous):

So I was correct